will like asap!
Question 36 of 59 1 Points Given that f(x) dx = -2, J-S f(x) dx = -1 and f(x) dx = 5 what is 380m) ds?
Answer #1

Given, I fead de=-2 il sex) dx = -1 7 seaddr = 1 c Na Sakota Series comment fair at senian 5 -3 [ 1 + (-2) + s] =3(1-2+5) =3(